# Multiple quantum exceptional, diabolical, and hybrid points in multimode bosonic systems: II. Nonconventional $\mathcal{PT}$-symmetric dynamics and unidirectional coupling

### v1

@@ V.1 Jordan canonical form of a general matrix
V.2 Perturbation of a dynamical matrix
The second approach is based upon introducing a suitable small perturbation of a dynamical matrix $\bm{M}$ that can remove both EDs and DDs. In this approach, the eigenvalues of the perturbed dynamical matrix $\bm{M_{\delta}}$ are determined and the degeneracies are removed as the degenerated eigenvalues split and then gradually diverge with the increasing perturbation. Perturbation can also be used for characterizing ED [32, 33, 34].
@@ V.2 Perturbation of a dynamical matrix
$F=\frac{\langle\bm{y_{2}}|\bm{y_{1}}\rangle}{\sqrt{\langle\bm{y_{1}}|\bm{y_{1}}\rangle\langle\bm{y_{2}}|\bm{y_{2}}\rangle}}$ (117)
and symbol $\langle$|$\rangle$ | ⟩ stands for the scalar product of complex vectors.
@@
    A two-mode system with usual coupling and $\beta=0$ described by the perturbed dynamical matrix
@@
    A two-mode system with usual coupling and $\beta=0$ described by the perturbed dynamical matrix
@@
    The eigenvalues and the corresponding eigenvectors are attained, respectively, in the form:
@@
(a) Refer to caption
(b) Refer to caption
(c) Refer to caption
@@
    A two-mode system with usual coupling and $\beta=0$ described by the dynamical matrix $\bm{M^{(2)}}$ where
@@
    $\bm{y^{\xi_{\delta,1}}_{1,2}}$ $=$ $\left[-\frac{\epsilon+\lambda^{\xi,\delta,1}_{1,2}}{\kappa},1\right]^{T}.$ (129)
@@
    A two-mode system with usual coupling and $\beta=0$ described by the dynamical matrix $\bm{M^{(2)}}$ where
@@
    $\bm{y^{\xi_{\delta,2}}_{1,2}}$ $=$ $\left[-\frac{\epsilon+\lambda^{\xi+\delta}_{1,2}}{\kappa},1\right]^{T}.$ (132)
@@
where $4\gamma_{\pm}=\gamma_{1}\pm\gamma_{2}$ and $\beta^{2}={{\xi}^{2}-\gamma_{-}^{2}}$. The real and imaginary parts of the eigenvalues $\Lambda_{j}$, $j=1,\ldots,4$, of the $4\times 4$ matrix $\bm{M^{(2)}}$ are plotted in Fig. LABEL:fig5.
(a) Refer to caption
(b) Refer to caption
@@
Finally, we mention a specific case of the perturbation $\delta$ that removes DD but keeps one from the two originally diabolically-degenerated QEPs present in the system. The dynamical matrix $\bm{M^{(2)}}$ of the two-mode bosonic system with usual coupling is perturbed in the following way:
@@
VI Higher-order hybrid points revealed by field-operator moments
Let us consider a bosonic system composed on $n$ modes and described by an appropriate quadratic non-Hermitian Hamiltonian. Such a system is described by the $2n$ annihilation and creation operators, and the dynamical matrix $\bm{M^{(n)}}$ of its Heisenberg-Langevin equations has $2n$ eigenvalues. Let us fix a position in the system parameter space. After diagonalization of the dynamical matrix $\bm{M^{(n)}}$, we identify the eigenvalues $\lambda_{j}$ with different eigenvectors and count their degeneration numbers $n_{j}$ according to the number of coalescing eigenvectors (from the original maximal Hilbert space). Denoting the number of such eigenvalues by $\Sigma_{\lambda}$, we have
$\sum_{j=1}^{\Sigma_{\lambda}}n_{j}=2n.$ (143)
We assign, to any eigenvalue $\lambda_{j}$, an operator vector $\bm{\hat{B}_{j}}$ that encompasses the diagonalized field operators $\hat{b}_{l}$ and $\hat{b}_{l}^{\dagger}$ ($l=1,\ldots,n$) that are associated with the coalescing vectors of this eigenvalue.
Now we consider the dynamics of $k$th-order FOMs. These $k$th-order FOMs can formally be expressed in their general form using the above-defined operator vectors $\bm{\hat{B}_{j}}$ as $\langle\prod_{j=1}^{\Sigma_{\lambda}}\bm{\hat{B}_{j}}^{k_{j}}\rangle$, where the nonnegative integers $k_{j}$ obey $\sum_{j=1}^{\Sigma_{\lambda}}k_{j}=k$. The vector $\bm{k}$ defined as $(k_{1},k_{2},\ldots,k_{\Sigma_{\lambda}})$ then points out at a possible QHP with the complex eigenfrequency $\lambda$ given as
$\lambda=\sum_{j=1}^{\Sigma_{\lambda}}\lambda_{j}^{k_{j}}.$ (144)
The degrees $d_{\rm e}^{\bm{k}}$ of ED and $d_{\rm d}^{\bm{k}}$ of DD of this QHP are given by the following formulas:
$d_{\rm e}^{\bm{k}}$ $=$ $\prod_{j=1}^{\Sigma_{\lambda}}n_{j}^{k_{j}},$ (145)
$d_{\rm d}^{\bm{k}}$ $=$ $\frac{k!}{\prod_{j=1}^{\Sigma_{\lambda}}k_{j}!}.$ (146)
The number $N^{(k)}_{B}$ of such QHPs observed for the $k$th-order FOMs (see the fourth columns of Tabs. I—VI in Ref. [8]) is determined using the combination number as:
$N^{(k)}_{B}=\left(\begin{array}[]{c}\Sigma_{\lambda}+k-1\\ k\end{array}\right).$ (147)
Similar formula also gives the number $N^{(k)}_{b}$ of different $k$th-order FOMs written in the operators $\hat{b}_{l}$ and $\hat{b}_{l}^{\dagger}$ for $l=1\ldots n$:
$N^{(k)}_{b}=\left(\begin{array}[]{c}2n+k-1\\ k\end{array}\right).$ (148)
These FOMs are explicitly written in the third columns of Tabs. I—VI in Ref. [8]. Also, the number $N^{(\bm{k})}_{b}$ of different $k$th-order FOMs written in the operators $\hat{b}_{l}$ and $\hat{b}_{l}^{\dagger}$ and belonging to a fixed $\bm{k}$ is expressed as:
$N^{(\bm{k})}_{b}=\prod_{l=1}^{k}\left(\begin{array}[]{c}n_{l}+k_{l}-1\\ k_{l}\end{array}\right).$ (149)
To demonstrate the general approach and formulas, we apply them to the four-mode bosonic system in the circular configuration analyzed in Sec. II in the regime of nonconventional PT-symmetric dynamics. Its first- and second-order FOMs and the revealed QHPs with their degeneracies are given in Tab. 1. The general approach results in Tab. 4 that allows to easily derive the content of Tab. 1: Two lines belonging to $\bm{\Lambda_{j}^{\rm i}}=\gamma_{+}$ are obtained assuming $\bm{k}=(0,0,0,0,1,0)$ and $(0,0,0,0,0,1)$. For $\bm{\Lambda_{j}^{\rm i}}=2\gamma_{+}$ the entries to Tab. 1 are in turn generated assuming $\bm{k}=(0,0,0,0,1,1)$, $(0,0,0,0,2,0)$, and $(0,0,0,0,0,2)$. The orders of the corresponding EDs and DDs are derived using Eqs. (145) and (146), respectively.
$\lambda_{j}$ $n_{j}$ $\bm{\hat{B}_{j}}$ $k_{j}$
$\gamma_{13}$ 1 $\bm{\hat{B}_{1}}\equiv[\hat{b}_{1}]$ $k_{1}=0$
$\gamma_{13}$ 1 $\bm{\hat{B}_{2}}\equiv[\hat{b}_{1}^{\dagger}]$ $k_{2}=0$
$\gamma_{24}$ 1 $\bm{\hat{B}_{3}}\equiv[\hat{b}_{2}]$ $k_{3}=0$
$\gamma_{24}$ 1 $\bm{\hat{B}_{4}}\equiv[\hat{b}_{2}^{\dagger}]$ $k_{4}=0$
$\gamma_{+}$ 2 $\bm{\hat{B}_{5}}\equiv[\hat{b}_{3},\hat{b}_{3}^{\dagger}]$ $k_{5}$
$\gamma_{+}$ 2 $\bm{\hat{B}_{6}}\equiv[\hat{b}_{4},\hat{b}_{4}^{\dagger}]$ $k_{6}$
Table 4: Eigenvalues $\lambda_{j}$, their degeneracies $n_{j}$, the corresponding field-operator vectors $\bm{\hat{B}_{j}}$, and their varying powers $k_{j}$ for the four-mode bosonic system analyzed in Tab. 1. QHPs are uniquely identified by powers of $k_{5}$ and $k_{6}$.
VII Conclusions
We have continued the analysis of the dynamics of simple bosonic systems described by quadratic non-Hermitian Hamiltonians from the point of view of the occurrence of quantum exceptional, diabolical, and hybrid points. We have revealed the bosonic systems exhibiting nonconventional $\mathcal{PT}$-symmetric dynamics characterized by the observation of quantum exceptional and hybrid points only in certain subspace(s) of the whole system Liouville space. We have identified such nonconventional second-order inherited quantum exceptional and hybrid points in four-mode bosonic systems.
Applying the method of concatenating simple bosonic systems via unidirectional coupling, we have found the conditions for the observation of up to sixth-order inherited quantum exceptional and hybrid points. However, by analyzing the behavior of the two-mode bosonic system with unidirectional coupling we have shown that the bosonic systems with unidirectional coupling are applicable only for short times in which they ensure physically consistent behavior. In short times, more complex bosonic systems with diverse structures and arbitrary-order exceptional degeneracies can be built by concatenating simple bosonic systems via unidirectional coupling of several types. Nevertheless, tailoring the properties of the Langevin operator stochastic forces in systems with unidirectional coupling does not enable extending their applicability to arbitrary times, in contrast to how the damping and amplification are consistently described.
Using analytical formulas for two-mode bosonic systems we have elucidated the operation of two numerical methods for the identification of quantum exceptional points and their degeneracies. These include: (1) the transformation of a dynamical matrix into its Jordan form and (2) the introduction of a suitable perturbation $\delta$ into the dynamical matrix and its subsequent eigenvalue analysis.
The exceptional and diabolical degeneracies of inherited quantum hybrid points have been used to derive higher-order degeneracies observed in the dynamics of higher-order field-operator moments. The quantum exceptional and hybrid points of second-order field-operator moments have been summarized in tables that evidence a rich dynamics of the field-operator moments.
Numbers of genuine and induced quantum hybrid points and their exceptional and diabolical degeneracies have been expressed as they depend on the order of the field-operator moments in the general form using parameters of the inherited quantum exceptional and hybrid points and their degeneracies.
This analysis considerably broadens the investigations of bosonic systems with $\mathcal{PT}$-symmetry exhibiting correct physical behavior, though it does not reveal bosonic systems with higher-order exceptional and hybrid singularities found for arbitrarily long times. The bosonic systems with unidirectional coupling enable generating higher-order exceptional and hybrid points only in their short-time dynamics. The systems with nonconventional $\mathcal{PT}$-symmetric dynamics exhibit only low- (second-) order exceptional and hybrid degeneracies, similarly as their counterparts analyzed in Ref. [8]. This analysis together with that in Ref. [8] show that the observation of higher-order exceptional and hybrid singularities in real, i.e. physically well-behaved, bosonic systems at arbitrary times is challenging.
VIII Acknowledgements
The authors thank Ievgen I. Arkhipov for useful discussions. J.P. and K.T. acknowledge support by the project OP JAC CZ.02.01.01/00/22_008/0004596 of the Ministry of Education, Youth, and Sports of the Czech Republic. A.K.-K., G.Ch., and A.M. were supported by the Polish National Science Centre (NCN) under the Maestro Grant No. DEC-2019/34/A/ST2/00081.
@@ VIII Acknowledgements
Appendix A Two-mode system with unidirectional coupling and relevant reservoir properties
@@
$\langle\hat{\bm{L^{\rm u}}}(t)\hat{\bm{L}^{{\rm u}\dagger}}(t^{\prime})\rangle$ $=$ $\delta(t-t^{\prime})$ (160)
×$\times\left[\begin{array}[]{cccc}2\gamma&0&-i\epsilon l_{1}(t)&0\\ 0&0&0&0\\ i\epsilon l_{1}(t)&0&-\frac{\epsilon^{2}\mu(t)^{2}l_{2}(t)}{\gamma}&\frac{\epsilon\kappa l_{1}(t)l_{2}(t)}{2\gamma}\\ 0&0&\frac{\epsilon\kappa l_{1}(t)l_{2}(t)}{2\gamma}&2\gamma-\frac{\kappa^{2}l_{2}(t)}{\gamma}\end{array}\right],$
@@
Appendix B Statistical properties of a two-mode bosonic system with unidirectional coupling
We consider modes 1 and 2 being in their initial coherent states $\alpha_{1}$ and $\alpha_{2}$, respectively. The modes evolution described in Eq. (LABEL:52) maintains the state Gaussian form described by the following normal characteristic function [11, 37, 7]:
$C_{\cal N}(\beta_{1},\beta_{2};t)$ $=$ $\exp\left[-B_{2}(t)|\beta_{2}|^{2}+\{C_{2}(t)\beta_{2}^{*2}+{\rm c.c.}\}/2\right]$
×$\times\exp[\alpha_{1}^{*}(t)\beta_{1}+\alpha_{2}^{*}(t)\beta_{2}-D(t)\beta_{1}^{*}\beta_{2}^{*}+{\rm c.c.}],$

### v2

@@ B.1 Jordan canonical form of a general matrix
B.2 Perturbation of a dynamical matrix
The second approach is based upon introducing a suitable small perturbation on a dynamical matrix $\bm{M}$ that can remove both EDs and DDs. In this approach, the eigenvalues of the perturbed dynamical matrix $\bm{M_{\delta}}$ are determined and the degeneracies are removed as the degenerated eigenvalues split and then gradually diverge with the increasing perturbation. Perturbation can also be used for characterizing ED [57, 58, 59].
@@ B.2 Perturbation of a dynamical matrix
$F=\frac{|\langle\bm{y_{2}}|\bm{y_{1}}\rangle|}{\sqrt{\langle\bm{y_{1}}|\bm{y_{1}}\rangle\langle\bm{y_{2}}|\bm{y_{2}}\rangle}}$ (144)
and symbol $\langle|\rangle$ stands for the scalar product of complex vectors.
@@
    A two-mode system with bidirectional coupling and $\beta=0$ described by the perturbed dynamical matrix
@@
    A two-mode system with bidirectional coupling and $\beta=0$ described by the perturbed dynamical matrix
@@
    The eigenvalues and the corresponding eigenvectors are obtained, respectively, in the form:
@@
(a) Refer to caption (b) Refer to caption (c) Refer to caption
@@
    A two-mode system with bidirectional coupling and $\beta=0$ described by the dynamical matrix $\bm{M^{(2)}}$ where
@@
    $\bm{y^{\xi_{\delta,1}}_{1,2}}$ $=$ $\left[-\frac{\epsilon+\lambda^{\xi_{\delta,1}}_{1,2}}{\kappa},1\right]^{T}.$ (156)
@@
    A two-mode system with bidirectinal coupling and $\beta=0$ described by the dynamical matrix $\bm{M^{(2)}}$ where
@@
    $\bm{y^{\xi_{\delta,2}}_{1,2}}$ $=$ $\left[-\frac{\epsilon+\lambda^{\xi_{\delta,2}+\delta}_{1,2}}{\kappa},1\right]^{T}.$ (159)
@@
where $4\gamma_{\pm}=\gamma_{1}\pm\gamma_{2}$ and $\beta^{2}={{\xi}^{2}-\gamma_{-}^{2}}$. The real and imaginary parts of the eigenvalues $\Lambda_{j}$, $j=1,\ldots,4$, of the $4\times 4$ matrix $\bm{M^{(2)}}$ are plotted in Fig. 6.
(a) Refer to caption (b) Refer to caption
@@
Finally, we mention a specific case of the perturbation $\delta$ that removes DD but keeps one from the two originally diabolically-degenerated QEPs present in the system. The dynamical matrix $\bm{M^{(2)}}$ of the two-mode bosonic system with bidirectional coupling is perturbed in the following way:
@@
Appendix C Two-mode bosonic system with unidirectional coupling and relevant reservoir properties
@@
$\langle\hat{\bm{L^{\rm u}}}(t)\hat{\bm{L}^{{\rm u}\dagger}}(t^{\prime})\rangle$ $=$ $\delta(t-t^{\prime})\left[\begin{array}[]{cccc}2\gamma&0&-i\epsilon l_{1}(t)&0\\ 0&0&0&0\\ i\epsilon l_{1}(t)&0&-\frac{\epsilon^{2}\mu(t)^{2}l_{2}(t)}{\gamma}&\frac{\epsilon\kappa l_{1}(t)l_{2}(t)}{2\gamma}\\ 0&0&\frac{\epsilon\kappa l_{1}(t)l_{2}(t)}{2\gamma}&2\gamma-\frac{\kappa^{2}l_{2}(t)}{\gamma}\end{array}\right],$ (180)
@@
Appendix D Statistical properties of a two-mode bosonic system with unidirectional coupling
We consider modes 1 and 2 being in their initial coherent states $\alpha_{1}$ and $\alpha_{2}$, respectively. The modes evolution described in Eq. (LABEL:54) maintains the state Gaussian form described by the following normal characteristic function [32, 56, 11]:
$C_{\cal N}(\beta_{1},\beta_{2};t)$ $=$ $\exp\left[-B_{2}(t)|\beta_{2}|^{2}+\{C_{2}(t)\beta_{2}^{*2}+{\rm c.c.}\}/2\right]$ (182)
$\hskip-42.67912pt\times\exp[\alpha_{1}^{*}(t)\beta_{1}+\alpha_{2}^{*}(t)\beta_{2}-D(t)\beta_{1}^{*}\beta_{2}^{*}+{\rm c.c.}],$